How many gallons of a 2% acid solution must be mixed with 7 gallons of a 20% solution to obtain a 17% solution?
Answer by ptaylor(2198)

How many gallons of a 2% acid solution must be mixed with 7 gallons of a 20% solution to obtain a 17% solution?
Let x=amt of 2% solution needed
Now we know that the amount of pure acid that exists before the mixture takes place (0.02x+0.20*7) has to equal the amount of pure acid that exists after the mixture takes place(0.17(7+x))
Sooooo
0.02x+0.20*7=0.17(7+x)
0.02x+1.4=1.19+0.17x
0.17x-0.02x=1.4-1.19
0.15x=0.21
x=1.4 gal---amount of 2% solution needed
CK
0.02(1.4)+1.4=0.17(8.4)
0.028+1.4=1.428
1.428=1.428
Hope this helps----ptaylor